If $2000.00 invested at 10% simple annual interest, how much should b invested at 12% from both so the yearly income is $5,000.00?
Answer by addingup(3677)

2000*.1 + .12(x-2000) note that the amount invested at 12% is x-2000.
Let's solve for x:
200+.12x-240 = 5000
.12x-40 = 5000
.12x = 5040
x = 42000 this is the total amount invested.
Now:
At 10% = 2000
At 12% = 42000-2000 = 40000
--------------------------------------
Check:
.10*2000 = 200
.12*40000 = 4800
200+4800 = 5000 Correct